A 0.100 M weak acid is titrated with a strong base to the equivalence point. The pH of

the resulting solution is found to be 9.18.

What is the pKa of the acid?
Answer #1

Please do give a thumbs up.
Answer #2

Calculation:

  1. Given:

    • pH at equivalence point = 9.18 (solution is basic).

    • Weak acid (HA) is fully neutralized to its conjugate base (A⁻) by the strong base.

  2. Find [OH⁻] from pH:

    pOH=14pH=149.18=4.82[OH]=10pOH=104.821.51×105 M

  3. Set up the hydrolysis reaction of A⁻:

    A+H2OHA+OH

    At equilibrium:

    Kb=[OH][HA][A](1.51×105)20.100=2.28×109

  4. Relate Kb to Ka:

    Ka×Kb=1×1014Ka=1014Kb=10142.28×1094.39×106

  5. Calculate pKa:

    pKa=log(Ka)=log(4.39×106)5.36


Conclusion: The pKa of the weak acid is 5.36.


answered by: Harshwardhan kunal